Back to the NIKA2 main page

Daily Reports

Feb 22nd, Tuesday

00:00-03:00; Albrecht, Angel: DAQ started at 00:15; bad KIDs 44/5/46 (ar1/ar2/ar3). Tau(225 GHz) < 0.2 throughout this mini-shift. FWHM 17.2" x 17.9" (2 mm); quite stable atmosphere: p2p variations of ~0.5 Jy/b, same band. @calib_1scan on CRL618 and ALFORI (scans 20220221s275 and s277). Then a @skydip (s278). Switch to project 192-16. Four rounds launched on COSMOS field (the last one running in the next shift).

03:00-09:00; Corentin: Finished the repetition on COSMOS. Then I switched to nikaw-21 project and did a pointing & focus and got a beammap on 1253-055. Then a skydip, there was a little increase in opacity (0.2 --> 0.3) but it decreased rapidly to 0.2. I switched to project 199-16, and did 2 reps on PSZ2G081 (32-->34/44) (with a point on 1633+382), then 4 reps on PSZ2G094 (20-->24/24 FINISHED)(with a point + focus on 1641+399). Finally a rep on PSZ2G081 (35/44) and switched to project nikaw-21 in order to do 3 @calib1scan on MWC349, NGC7027, and CRL2688 (20220222s89, s90, and s91). Finishing with a @skydip (s92). Opacity remained stable quite all night (~0.2).

09:00-11:50; Angel: Starting on project 109-21. Two reps. (=4 rounds) on B68. Got @calib_1scan on Mars (s117). Then, 1 rep. on B335. Opacity was below 0.2. Nicely stable atmosphere.

15:40-21:00; Angel: Checking system after maintenance. Computer Mac8 doesn't work after several tries. Retuning; bad KIDs 43/4/47 (ar1/ar2/ar3). Opacity < 0.2. Point & focus on 0316+413, looking for starting with project 122-16 (GASTON). Imbfits creation takes unexpectedly long times; mrt-lx3 is slow. Reported to the IT group. NFS service anomalies, with nika2-a and mrt-lx3 involved, were detected and restored by PM & WR. Consequently, DAQ was restarted; retuning, bad KIDs 31/5/43 (ar1/ar2/ar3). Checking pointing & focus on CRL618. The system is stable again. GASTON targets were too high at that time. Got a @calib_1scan on CRL618 (s152) and ALFORI (s155). Switching to project 143-21; quite stable atmosphere. Target G202 observed: 1 rep. Focus and pointing on Uranus + @calib_1scan (s182) + @skydip (s183). Beams are not in good shape for a beam map. Asking SS for checking the back structure/main dish temperatures. Doing the overlap with Matthieu.

21:00-03:00; Matthieu: We started the shift with focus on Uranus followed by 0316+413. The beam seemed finally stable after the difficulty of the afternoon and we did a beammap [s202]. We then did a pointing and a calib_1scan on CRL618. We then observed GASTON [l1521e] during ~1h and finally moved to COSMOS until the rest of the shift. The weather was stable with a tau going from 0.2 in the early night to 0.15.

Feb 23rd, Wednesday

03:00-09:00; Karine, Isabelle: Following the previous shift, we continued observing project 192-16. We did a focus that we had to repeat once to get reasonable results and nice beams. Then we did 2 reps (cosmos). At 4h30 the wind was too high and the telescope was parked until 7h20. We resume observations at 7h40 [retuning: bad KIDs 42/5/46 (ar1/ar2/ar3)] with a pointing/focus on 1757-240 on nikaw-21 account. Then we did one rep on B68 (project 109-21). Tau was ~0.2 at the beginning of the shift and increased to 0.26 by the end of the shift.

09:00-15:00; Andrew: Going to MWC349 for standard calibrations. Stopped at 09:05 because of wind gusts. Resuming observations with NIKA2 at 11:40, but with not-so-good perspectives. Retuning: bad KIDs 53/6/53 (ar1/ar2/ar3). Pointing to 2251+158 for getting focus and monitoring stability. Good beam sizes. A @calib_1scan on Mars (20220223s61) and @skydip (s62). Opacity increasing above 0.35 and unstable. Switched to EMIR (~13:30) backup program 174-21 for getting 0.5 reps. on 23028 before tau(225 GHz) > 0.7. Telescope stopped at 14:45.

15:00-21:00; Jean-Francois (on-hold), Angel: Stopped until ~18:45 when tau(225 GHz) dropped to ~0.4, unstable. Restarting observations with EMIR, backup project 116-21 (E0 band). Target L1544 observed (8 reps. of 7 min. each). Preparing overlap for Hélène and Frédéric.

21:00-03:00; Hélène, Frédéric: We conducted only EMIR observations in band E090 at 83 Ghz for program 116-21 (source L1544). Including the repetitions done in the previous shift, the accumulated time is 16 repetitions * 12 min ~ 3.2 h. No incident to report. The zenith opacity fluctuated wildly, with smoothed averages of ~ 0.5 to 0.3. The focus went from -1.75 to ~ -1.6 and back to -1.75. After the last repetition, the source declination was about 25 degrees, too low to continue. As no other program could be observed in such conditions, we stopped and the antenna was parked at about 1 am. [AB: Too unstable to continue on project 170-21 (EMIR E0+E2 bands)].

Feb 24th, Thursday

03:00-09:00; Karine, Roberta (on-hold): [Cloudy. Trying to resume observations with a first rising target of EMIR project 174-21 (E0 band) at 07:45. Opacity increasing above 0.7, large variations. Telescope stopped.]

09:00-15:00; Andrew (on-hold): [Telescope stopped. A brief snowfall, wind gusts, <tau> > 0.6 with wild variations].

15:00-21:00; Aurore Bacmann (on-hold): [tau < 0.5 at the beginning. Trying project 174-21 (21272) with EMIR (E0). Ugly pointing corrections (anomalous refraction)/poor focus fittings. PWV > 12 mm. Telescope stopped. Dense fog, snowfall. Deicing switched on at 19:00].

21:00-03:00; Alexandre (on-hold): [Fog, snow]

Feb 25th, Friday

03:00-09:00; Frédéric (on-hold): [Stop snowing; high wind speed. Deicing switched off at 07:00.]

09:00-15:00; Xavier (on-hold): [Sky cleared for about 2 hours, but unstable. Thick fog again starting ~11h. Sparsed rain/snow with son gaps in the middle. Unstable.]

15:00-21:00; Jean-Francois (on-hold): [Mean opacity between 0.4 and 0.6. Restarting observations; we tried with EMIR's 116-21 project at ~17:30. Rain & wind alarms triggered after a while. Telescope parked.]

21:00-03:00; Hélène, Frédéric (on-hold): [Opacity about 0.6 at the beginning of the shift; wind gusts with speed > 20 m/s, light snow.]

Feb 26th, Saturday

03:00-09:00; Emmanuel: [Mild winds, snowing stopped, opacity tends to go down; deicing switched on between 3h and ~5h, main dish only.] Observations started with EMIR -the forecasts are not so promising for this morning- for project 170-21 (E0 & E2 bands), but trying first under project test. Looks acceptable at 3 mm, but not so at 1 mm. Switched to EMIR project 174-21 (E0 band only). Target 20286 observed, 1 repetition.

09:00-15:00; Juan: Continuing of project 174-21. Some instabilities prevented good focus/pointing corrections at the beginning, tau ~0.5. Got 5 out of 8 scans on target 20286 before stopping the telescope at 10:05 because of high wind speed.

15-00-21:00; Quang (on-hold): [Stopped. High wind speed, relative humidity of 100%].

21:00-03:00; Nicolas (on-hold): [Fog. No wind.]

Feb 27th, Sunday

03:00-09:00; Antonio, Marco (on-hold): [Fog. Light snowfall at ~sunrise time].

09:00-15:00; Angel: [Thick Fog. Light, sparsed snowfalls continued. At ~14:30 we had a power failure in the cooling system of NIKA2 Helium compressors. Restarted after several tries: OK. The Receiver Group is working on the eventual cooling down. Observations with NIKA2 will be feasible when they give us the green light.]

15:00-21:00; Jean-Francois (on-hold): [Thick fog. Occasional light rainfalls.]

21:00-03:00; Hélène: Started at ~10 pm with EMIR project 116-21. Did in total 6 more repetitions on source L1544. From 11:10 pm on, the baselines became strongly non-linear. Around midnight, it was time for a new focus, but it was impossible to point even on 0355+508, because of the fog. We decided to switch to project 170-21 a bit earlier than planned, but could do nothing, because the fog continued and the snow forced us to stop at 1 am. There was 100% humidity for the whole shift, and the 225 GHz opacity fluctuated a lot. [Deicing connected for ~30' at 02:15.]

Feb 28th, Monday

03:00-09:00; Aishwarya (on-hold for NIKA2), Mónica, Angel (overlapping after 07h): Resuming observations with EMIR backup project 170-21. Targets 41455-high and 51442-high were observed. Atmospheric instabilities and fog after ~06:30. Switched to EMIR project 174-21 (E0 band only). Hard to get pointing/focus corrections. 100% humidity. Thick fog. Stopped at 07:15, waiting for better conditions.

09:00-15:00; Xavier: Observations with EMIR resumed at ~10:15. Target 23129 observed (1 rep.; finished), then 1 rep. on target 20286. Becoming unstable at ~13:30. Switched to project 'test'. Signs of anomalous refraction. Hard to get pointing/focus corrections. Telescope parked at ~14:45.

15:00-21:00; Aurore: [Still unstable, despite mean tau drops. From the Receiver Group (16:30, many thanks for the quick fix), NIKA2 is ready again for science operations. DAQ restarted. Atmosphere instabilities persist. Light snow for a while. Retuning on-sky, bad KIDs (ar1/ar2/ar3) 39/5/48. Checking noise traces: OK.] NIKA2 observations start again [17:45], with low opacity and stability improving while the Sun goes down. Performed [a calib_1scan on Uranus, s121] a skydip (Az Uranus) [s122] and a calib1scan on ALFORI [s125]. Then switched to project 143-21 for one repeat. At ~ 20:30, switched to 122-16 for one repeat, but the wind increased. Beams were fine (typically 17"x18" at 2mm).

21:00-03:00; Guilaine: Finished the scans on L1521e and launched one more repetition that was aborted at 21:14 due to high wind. Waiting for about one hour to be able to start observing again (tau=0.08). Went to QSO 0316+413, pointing/focus/beammap [s153]. Went to CRL618: pointing/calib_1scan [s155]. Did a skydip [s156]. Went to 1055+018 pointing/focus. 23h30: started to observe COSMOS. Up to 3am, did 7 repets. Wonderful: didn’t see such stable and good beams for a long time.

Mar 1st, Tuesday

03:00-10:30; Juan: Very good weather, starting with focus and pointing after COSMOS scans (good beams); 2 repetitions on PSZ2G183.9; 4.5 repetitions in PSZ2G081; Switched to project 109-21, skydip [scan s60], 4 x 2 scans [= 2 reps.] in B68; Focus & Calibration on MWC349, NGC7027 and CRL2688 (2 x @calib_1scan; s82 to s87), 1.5 reps in PSZ2G046 (stopped because of a problem with mrt-lx1) [scan 96 got stuck].

DailyReportsNika2Pool220222 (last edited 2022-03-07 08:22:20 by NikaBolometer)